4. Five applicants (Jim, Don, Mary, Sue, and Nancy) are available for two identical jobs. Because all applicants are equally qualified, a supervisor randomly selects two applicants to fill these jobs. Find the probability of each of the following events a Both males are selected. b At least one male is selected c At least one female is selected.
Answer #1

Number of applicants = 5

Number of males = 2

Job Positions = 2

a) Probability of selecting both males =

2.4 = 0.1

b) Probability of selecting atleast one male = P(one male) +P(two male)

23 2 *1 0.3 0.1 = 0.4 5 4 5 *4

b) Probability of selecting atleast one female = P(one female) +P(two female)

3 *2 3 *2 545k4 0.3+ 0.30.6